OpenStudy (anonymous):

someone please help with algebra 1!!!!

OpenStudy (anonymous):

Okay?

OpenStudy (candy13106):

wats your question

OpenStudy (anonymous):

Iris has been studying an invasive population of snails. This particular snail has no local predators so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years. Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences. A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then explain to Iris how the key features of this local snail population compares to the key features of the invasive population. Iris wants to graph the invasive snail population to show the city council. Justify what the appropriate domain and range would be for the function f(x), what the y-intercept would be, and if the function is increasing or decreasing. In five years, a garden festival plans on using the park where Iris has been studying the invasive snails. Explain to the garden festival committee how to find the average rate of change for the snail population between years 2 and 5. Describe what this average rate of change represents.

OpenStudy (anonymous):

?? please help

OpenStudy (anonymous):

somebody? for a medal??

OpenStudy (candy13106):

idk sorry

OpenStudy (anonymous):

could you tag someon who can?

OpenStudy (candy13106):

idk who

OpenStudy (anonymous):

could someone please help?

OpenStudy (anonymous):

@ranga is the man for you.

OpenStudy (anonymous):

@ranga please help?

OpenStudy (anonymous):

??? @ranga

OpenStudy (ranga):

Exponential growth function looks like: y = A * (1+r)^x Here, A would be the initial population and r would the rate of growth. You can pick some numbers. Let there be 100 snails to start with and let the rate of growth be 5%. Then A = 100 and r = 0.05. Plug it into the equation: y = 100 * (1+0.05)^x y = 100(1.05)^x where x is the time in years and y is the snail population.

OpenStudy (anonymous):

thats for the first problem?

OpenStudy (anonymous):

how would i do the next one?

OpenStudy (ranga):

For the first problem also include that domain of the function is [0,15] because they say it is an exponential growth for 15 years and so the formula is good only for the first 15 years.

OpenStudy (anonymous):

okay i will.

OpenStudy (ranga):

\[\Large g(x) = 200(1.03)^{2x} = 200((1.03)^2)^x = 200(1.0609)^x\]

OpenStudy (ranga):

We have converted the function with exponent 2x to just x above. g(x) = 200(1 + 0.0609)^x Compare this to the general function stated in the first part: y = A(1+r)^x We find A = 200 and r = 0.0609. Therefore the initial snail population here is 200 and the rate of growth is 0.0609 or 6.09%

OpenStudy (anonymous):

okay so the third one?

OpenStudy (ranga):

From part 1) we have f(x) = 100(1.05)^x when x = 0, f(x) = 100(1.05)^0 = 100. This is the y-intercept which represents the starting population. Every year the population increases by 5% and therefore this function will always be increasing. The domain is x = 0 to x = 15 because it is given in the problem: "She has observed that the population follows an exponential rate of growth for fifteen years." Domain is [0, 15] To get the range put x = 0 and x = 15 and calculate f(x). f(0) = 100(1.05)^0 = 100 f(15) = 100(1.05)^15 = 207.89 rounded down to 207 because number of snails has to be a whole number. So range is [100, 207] 

OpenStudy (ranga):

For the last part, the average rate of change between years x = 2 and x = 5 is: { f(5) - f(2) } / { 5 - 2 } = ? I will let you calculate the average rate using the above formula. What the average rate here represents would be the average number of increase in snail population each year between the years 2 and 5.

OpenStudy (anonymous):

so thats all the problems then? thats all i have to do? seems simple enough! thank you so much!

OpenStudy (ranga):

You are welcome.

OpenStudy (anonymous):

for the last one would the vaerage rate of change be 3/3

OpenStudy (anonymous):

??

OpenStudy (ranga):

Not 3/3. f(x) = 100(1.05)^x f(5) = 100(1.05)^5 = ? f(2) = 100(1.05)^2 = ? Plug the numbers into the formula: Average rate of change between years x = 2 and x = 5 is: { f(5) - f(2) } / { 5 - 2 }

OpenStudy (anonymous):

127.628-110.25/3?

OpenStudy (ranga):

Yes. But simplify the number and round it.